OpenStudy (anonymous):

Given the trinomial 2x*2 + 4x − 2, what is the value of the discriminant? a. 0 b. 10 c. 32 d. 36

OpenStudy (anonymous):

2x2+4x−2 comparing it with ax2+bx+c we find a= 2, b= 4 & c=2 now discriminant is given as: b2−4ac=42−4××2×2=16−16=0 so A.0 is the correct answer.

OpenStudy (anonymous):

C is the answer.. 32

OpenStudy (anonymous):

:3

OpenStudy (anonymous):